rewrite the equation in function form:

-3x-4y=6


3/2(x-6y)=-8

Answer by algebrapro18(206)

that must means solve for y

-3x-4y=6
-4y=3x+6
y = -3/4x - 3/2

3/2(x-6y)=-8
x-6y=-16/3
-6y=x-16/3
y= -1/6x - 8/9